### v2

@@ 6 Conclusions
We see several avenues worth exploring and left for future work. First, whereas the rational expectations solution does not satisfy consistency, Nicolò et al. (2023) show that existence is guaranteed in dynamic, one-to-one, one-sided matching markets. Because the same externality is present in both two-sided and one-sided dynamic matching markets, understanding what sets apart these two environments is useful.
Second, whereas we propose two new solution concepts for dynamic matching markets, these are by no means the only solution concepts that satisfy consistency. Further understanding the set of solution concepts that satisfy consistency would allow us to delineate the limits of the predictions afforded by such solution concepts. Nicolò et al. (2024) is a step in this direction.
Finally, because important dynamic matching applications, like sequential assignment in school choice, involve many-to-one matching, extending our results to many-to-one matching markets is natural. As the work of Altinok (2019) and Liu (2023) highlights, new issues arise in the study of such markets and understanding what the analogue of consistency is in these markets is worth exploring.
@@ Appendix A Omitted statements and proofs from Section 4
For completeness, we collect in this section the definitions analogous to those in Section 4 that allow us to describe the construction of the candidate matchings and state Theorem 4.1 solely in terms of the model primitives.
@@
Given an economy $E^{T}$, a period $t$, and a matching through period $t-1$, $m^{t-1}$. The one-period economy induced by the collection $\{\varphi_{t}(m^{t-1},k):k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$ is the length-1 economy defined as follows:
@@
    Preferences $\{u_{\varphi_{t}}(a,\cdot):\pazocal{B}_{t}(m^{t-1})\cup\{a\}\mapsto\mathbb{R}:a\in\pazocal{A}_{t}(m^{t-1})\},\{v_{\varphi_{t}}(\cdot,b):\pazocal{A}_{t}(m^{t-1})\cup\{b\}\mapsto\mathbb{R}:b\in\pazocal{B}_{t}(m^{t-1})\}$ are given by:
@@
      For each agent $a\in\pazocal{A}_{t}(m^{t-1})$, $u_{\varphi_{t}}(a,b)\geq u_{\varphi_{t}}(a,a)$ only if $u(a,b)\geq\min\{U_{t}(a,\overline{m}):\overline{m}\in\varphi_{t}(m^{t-1},a)\}$. Moreover, for $b,b^{\prime}\in\pazocal{B}_{t}(m^{t-1})$, $u_{\varphi_{t}}(a,b)\geq u_{\varphi_{t}}(a,b^{\prime})$ only if $u(a,b)\geq u(a,b^{\prime})$.
@@
      For each agent $b\in\pazocal{B}_{t}(m^{t-1})$, $v_{\varphi_{t}}(a,b)\geq v_{\varphi_{t}}(b,b)$ only if $v(a,b)\geq\min\{V_{t}(b,\overline{m}):\overline{m}\in\varphi_{t}(m^{t-1},b)\}$. Moreover, for $a,a^{\prime}\in\pazocal{A}_{t}(m^{t-1})$, $v_{\varphi_{t}}(a,b)\geq v_{\varphi_{t}}(a^{\prime},b)$ only if $v(a,b)\geq v(a^{\prime},b)$.
@@
where in a slight abuse of notation we use the notation $M^{\star}(\Phi)$ to signify the candidate matchings that can be constructed solely as a function of the family of conjectures, $\Phi$, for economy $E^{T}$.
@@
Fix $E^{T}$ and a matching $m^{\star}\in M^{\star}(\Phi)$. Suppose that for all $t\geq 1$ and all $k$ such that $m^{\star}_{t}(k)=k$, we have that
$m^{\star}\in\varphi_{t}(m^{\star^{t-1}},k).$ (A.1)
Then, for all $t\geq 1$, $(m_{s}^{\star})_{s=t}^{T}|_{E^{T}_{t}(m^{\star^{t-1}})}\in\Sigma_{T-t+1}^{\Phi}(E^{T}_{t}(m^{\star^{t-1}}))$.
@@
Assume that (A.1) holds. Note that for $t=T$ the result is vacuously true. Suppose we have established the result for $s\in\{t+1,\dots,T\}$. We show it holds for $s=t$. Note that $m^{\star}_{t}$ immediately satisfies parts 1 and 2 of 6 whenever $m^{\star}_{t}(k)\neq k$. Similarly, it satisfies part 3 of 6 for pairs $(a,b)\in\pazocal{A}_{t}(m^{t-1})\times\pazocal{B}_{t}(m^{t-1})$ such that $m^{\star}_{t}(a)\neq a,m^{\star}_{t}(b)\neq b$.
Consider now $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ such that $m^{\star}_{t}(k)=k$ and note that (A.1) implies that under $m^{\star}$, $k$ is getting at least the payoff of $k$’s worst conjectured matching under $\varphi_{t}(m^{\star^{t-1}},\cdot)$.
Finally, suppose $m^{\star}_{t}$ fails part 3 of 6 for some pair $(a,b)$ in $\pazocal{A}_{t}(m^{\star^{t-1}})\times\pazocal{B}_{t}(m^{\star^{t-1}})$. Then at least one of the members of the pair must be single under $m^{\star}_{t}$. Without loss of generality, assume $m^{\star}_{t}(a)=a$. Then, $u(a,b)>U_{t}(a,m^{\star})\geq\min\{U_{t}(a,\overline{m}):\overline{m}\in\varphi_{t}(m^{\star^{t-1}},a)\}$, where the second inequality follows from $m^{\star}\in\varphi_{t}(m^{\star^{t-1}},a)$. This chain of inequalities contradicts the definition of $m^{\star}_{t}$. Thus $(m_{s}^{\star})_{s=t}^{T}\in\Sigma_{T-t+1}^{\Phi}(E^{T}_{t}(m^{\star^{t-1}}))$ and the result follows. ∎
@@
In Section 4, we argued that the period-1 conjectures defined in Equation 4.1 lead to nonempty solutions for $E^{T}$ provided that the solution concept for the continuation economy, $\Sigma_{T-1}^{\Phi}$ is also nonempty. For completeness, we now build recursively the conjectures and the solution concept to complete the argument: A solution concept as in 6 built off of conjectures that rule out matchings that are not consistent with the solution concept in the continuation is always nonempty. In what follows, we keep the notation $\varphi^{*}$ for the conjectures and denote by $\Sigma^{\Phi^{*}}$ the induced solution concept.
Fix an economy of length $T$. Recall that 1 implies that no matter what the conjectures are, the set of $\Phi$-solutions for an economy of length $1$ coincide with the set of stable matchings. Given this, fix a matching through period $T-2$, $m^{T-2}$. For $k\in\pazocal{A}_{T-1}\left(m^{T-2}\right)\cup\pazocal{B}_{T-1}\left(m^{T-2}\right)$, define $k$’s conjectures as follows:
$\varphi^{*}_{T-1}(m^{T-2},k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{T-2},k):\overline{m}_{T}\in\mathrm{S}\left(E^{T}_{T}(\overline{m}^{T-1})\right)\right\}.$
Let $\Sigma^{\Phi^{*}}_{T-1}\left(E^{T}_{T-1}\left(m^{T-2}\right)\right)$ denote the set of $\varphi^{*}_{T-1}$-solutions for $E^{T}_{T-1}\left(m^{T-2}\right)$.
Recursively, suppose we have defined $\Sigma^{\Phi^{*}}_{T-t}$ for economies of length $T-t$, and fix a matching through period $t$, $m^{t-1}$. For $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, $k$’s conjectures are given by:
$\varphi^{*}_{t}(m^{t-1},k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1},k):(\overline{m}_{s})_{s=t+1}^{T}\in\Sigma^{\Phi^{*}}_{T-t}\left(E^{T}_{t+1}(\overline{m}^{t})\right)\right\},$
and let $\Sigma^{\Phi^{*}}_{T-t+1}\left(E^{T}_{t}(m^{t-1})\right)$ denote the set of solutions given the family $\Phi^{*}_{\geq t}=\{\varphi^{*}_{s}((m^{t-1},\overline{m}_{t}^{s-1}),k):s\geq t,\overline{m}_{t}^{s-1},k\in\pazocal{A}_{s}(m^{t-1},m_{t}^{s-1})\cup\pazocal{B}_{s}(m^{t-1},m_{t}^{s-1})\}$.
The following statement is an immediate corollary of Theorem 4.1 and the definition of the conjectures $\varphi^{*}$:
@@
For all $T\in\mathbb{N}$, $\Sigma^{\Phi^{*}}_{T}$ is a nonempty-valued correspondence.
A.1 then shows that existence issues for solution concepts in dynamic matching markets do not arise solely from the perfection requirement–i.e., that the continuation matching also abides by the solution concept–but from restrictions the conjectures place on what period-$t$ matchings are feasible.
@@
In what follows, we provide supporting details for the derivation in 4.1. First, note that should no one match in period $1$, a unique stable matching exists in $t=2$, which leads to the matching:
@@
Second, consider $E_{\setminus\{b_{1}\}}^{2}$. We argue that $m_{b_{1}}$ is a rational expectations solution. To see this, note that $b_{2}$ does not wish to block with $a_{1}$, and $a_{3}$ cannot do better than by matching in $t=1$ with $b_{2}$. The only possibility is that $b_{2}$ waits to be matched. However, $m_{\emptyset}$ is a valid conjecture for $b_{2}$ in $E_{\setminus\{b_{1},b_{2}\}}^{2}$, and $b_{2}$ prefers $a_{3}$ to matching with $a_{1}$ in $t=2$.
Third, consider $E_{\setminus\{a_{3},b_{1}\}}^{2}$. We show that $b_{1}$ is always matched to $a_{4}$ in any rational expectations solution and hence $b_{1}$ cannot match with $a_{1}$ in a rational expectations solution for $E_{\setminus\{a_{3}\}}^{2}$. In $E_{\setminus\{a_{3},b_{1}\}}^{2}$, three possibilities exist for a matching that is a rational expectations solution: (i) no one matches in $t=1$, which would lead to $m_{\emptyset}$ (though note that $m_{\emptyset}$ is blocked by $(a_{1},b_{2})$ in $E_{\setminus\{a_{3},b_{1}\}}^{2}$), (ii) $a_{1}$ matches with $b_{2}$, and (iii) $a_{2}$ matches with $b_{2}$. Under (ii), note that the outcome of side-$A$ proposing deferred acceptance in $t=2$ among $a_{2},a_{3},a_{4},b_{1},b_{3},b_{4}$ is
@@
so that any rational expectations solution for $E_{\setminus\{a_{3},b_{1}\}}^{2}$ that matches $a_{1},b_{2}$ in $t=1$ (if any) gives $b_{1}$ at least the payoff of matching with $a_{4}$. Similarly, under (iii), note that the outcome of side-$A$ proposing deferred acceptance in $t=2$ among $a_{1},a_{3},a_{4},b_{1},b_{3},b_{4}$ is
@@
so that any rational expectations solution for $E_{\setminus\{a_{3},b_{1}\}}^{2}$ that matches $a_{2},b_{2}$ in $t=1$ (if any) gives $b_{1}$ at least the payoff of matching with $b_{4}$. It follows that in $E_{\setminus\{a_{3}\}}^{2}$, $b_{1}$ has to match with $a_{4}$.
Fourth, consider $E_{\setminus\{a_{2}\}}^{2}$. We argue that $m_{a_{2}}$ is a rational expectations solution for that economy. To see this, we need to argue that $b_{1}$ cannot block. Consider then the economy $E_{\setminus\{a_{2},b_{1}\}}^{2}$ and note that $m_{b_{1}}$ is a rational expectations solution for the same reason it is a rational expectations solution in $E_{\setminus\{b_{1}\}}^{2}$. Thus, by blocking $m_{a_{2}}$, $b_{1}$ cannot expect to get more than the payoff of matching with $a_{1}$ in $t=2$. It follows that $m_{a_{2}}$ is a rational expectations solution for $E_{\setminus\{a_{2}\}}^{2}$.
Finally, we argue that $m^{\star}$ in the example is the outcome of the procedure discussed in Section 4. Note that because of $m_{a_{2}}$, $a_{2}$ is willing to match with $b_{2}$ in the one-period economy induced by the conjectures, and similarly, $b_{1}$ is willing to match with $a_{1}$. Because $a_{2}$ and $b_{2}$ top-rank each other in the one-period economy induced by the conjectures, they match together in any stable matching, which then justifies $a_{1}$ and $b_{1}$ matching, and $a_{3}$ remaining unmatched. When $\{a_{3},a_{4}\}$ are the $t=2$-agents on side $A$, a unique stable matching exists, which is that specified by $m^{\star}_{2}$.

### v3

@@ 6 Conclusions
We see several avenues worth exploring and left for future work. First, whereas the rational expectations solution does not satisfy consistency, Nicolò et al. (2023) show its existence is guaranteed in dynamic, one-to-one, one-sided matching markets. Because the same externality is present in both two-sided and one-sided dynamic matching markets, understanding what sets apart these two environments is useful.
Second, whereas we propose two new solution concepts for dynamic matching markets, they are by no means the only solution concepts that satisfy consistency. Further understanding the set of solution concepts that satisfy consistency would allow us to delineate the limits of the predictions afforded by such solution concepts. Nicolò et al. (2024) is a step in this direction. Relatedly, whereas consistent conjectures ensure the existence of $\Phi$-solutions, whether they are necessary for existence is worth exploring.
Finally, because important dynamic matching applications, such as sequential assignment in school choice, involve many-to-one matching, extending our results to many-to-one matching markets is natural. As Altinok (2019) and Liu (2023) highlight, new issues arise in the study of such markets, and understanding what the analogue of consistency is in these markets is worth exploring.
@@ Appendix A Omitted statements and proofs from Section 4
For completeness, in this section, we collect the definitions analogous to those in Section 4 that allow us to describe the construction of the candidate matchings and state the analogue of Theorem 4.1, Theorem A.1, solely in terms of the model primitives. The proof of Theorem 4.1 follows from that of Theorem A.1.
@@
Given an economy $E^{T}$, a period $t$, and a matching through period $t-1$, $m^{t-1}$, the one-period economy induced by the collection $\{\varphi_{t}(m^{t-1},k):k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$ is the length-1 economy defined as follows:
@@
    Utilities $\left\{u_{\varphi_{t}}(a,\cdot):\pazocal{B}_{t}(m^{t-1})\cup\{a\}\rightarrow\mathbb{R}|a\in\pazocal{A}_{t}(m^{t-1})\right\},\left\{v_{\varphi_{t}}(\cdot,b):\pazocal{A}_{t}(m^{t-1})\cup\{b\}\rightarrow\mathbb{R}|b\in\pazocal{B}_{t}(m^{t-1})\right\}$ are given by:
@@
      For each agent $a\in\pazocal{A}_{t}(m^{t-1})$, $u_{\varphi_{t}}(a,a)=\min\{U_{t}(a,\overline{m}):\overline{m}\in\varphi_{t}(m^{t-1},a)\}$. Moreover, for $b\in\pazocal{B}_{t}(m^{t-1})$, $u_{\varphi_{t}}(a,b)=u(a,b)$.
@@
      For each agent $b\in\pazocal{B}_{t}(m^{t-1})$, $v_{\varphi_{t}}(b,b)=\min\{V_{t}(b,\overline{m}):\overline{m}\in\varphi_{t}(m^{t-1},b)\}$. Moreover, for $a\in\pazocal{A}_{t}(m^{t-1})$, $v_{\varphi_{t}}(a,b)=v(a,b)$.
@@
where (i) in a slight abuse of notation, we use the notation $M^{\star}(\Phi)$ to signify the candidate matchings that can be constructed solely as a function of the family of conjectures, $\Phi$, for economy $E^{T}$, and (ii) $m^{\star}_{t}|_{E_{t,\varphi_{t}}(m^{\star^{t-1}})}$ is the restriction of $m^{\star}_{t}$ to the one-period economy $E_{t,\varphi_{t}}(m^{\star^{t-1}})$.
Definition 14.
Say $E^{T}$ has consistent conjectures if for all candidate matchings $m^{\star}\in M^{\star}(\Phi)$ the following holds:
For all t1 and all k such that mt(k)=k, we have that $\text{For all $t\geq 1$ and all $k$ such that $m^{\star}_{t}(k)=k$, we have that }m^{\star}\in\varphi_{t}(m^{\star^{t-1}},k).$ (A.1)
@@
Fix $E^{T}$ and suppose it has consistent conjectures. Then, the set of $\Phi$-solutions for $E^{T}$ is nonempty. In particular, $M^{\star}(\Phi)\subset\Sigma_{T}^{\Phi}(E^{T})$.
@@
Fix a candidate matching $m^{\star}\in M^{\star}(\Phi)$ that satisfies (A.1). Note that for all $t\in\{1,\dots,T\}$ and all $k\in\pazocal{A}_{t}(m^{\star^{t-1}})\cup\pazocal{B}_{t}(m^{\star^{t-1}})$, $m^{\star}$ satisfies parts 1 and 2 of 6 whenever $m^{\star}_{t}(k)\neq k$. Similarly, it satisfies part 3 of 6 for pairs $(a,b)\in\pazocal{A}_{t}(m^{t-1})\times\pazocal{B}_{t}(m^{t-1})$ such that $m^{\star}_{t}(a)\neq a,m^{\star}_{t}(b)\neq b$.
Consider now a period $t$ and an agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ such that $m^{\star}_{t}(k)=k$ and note (A.1) implies that under $m^{\star}$, $k$ is getting at least the payoff of $k$’s worst conjectured matching under $\varphi_{t}(m^{\star^{t-1}},k)$. Given the statement above, $m^{\star}$ satisfies parts 1 and 2 of 6 for all $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$.
Finally, suppose $m^{\star}_{t}$ fails part 3 of 6 for some pair $(a,b)$ in $\pazocal{A}_{t}(m^{\star^{t-1}})\times\pazocal{B}_{t}(m^{\star^{t-1}})$. Then at least one of the members of the pair must be single under $m^{\star}_{t}$. Without loss of generality, assume $m^{\star}_{t}(a)=a$. Then, $u(a,b)>U_{t}(a,m^{\star})\geq\min\{U_{t}(a,\overline{m}):\overline{m}\in\varphi_{t}(m^{\star^{t-1}},a)\}$, where the second inequality follows from $m^{\star}\in\varphi_{t}(m^{\star^{t-1}},a)$. This chain of inequalities contradicts the definition of $m^{\star}_{t}$. Thus, $m^{\star}\in\Sigma_{T}^{\Phi}(E^{T})$ and the result follows. ∎
@@
In Section 4, we argued that the period-1 conjectures defined in Equation 4.1 lead to nonempty solutions for $E^{T}$ provided the solution concept for the continuation economy, $\Sigma_{T-1}^{\Phi}$, is also nonempty. For completeness, we now define the conjectures and the solution concept determined by recursively applying Equation 4.1. This allows us to complete the argument implicit in that discussion. Specifically, a solution concept as defined in 6, with conjectures consisting of all matchings whose continuations are solutions in the continuation economy, is always nonempty. In what follows, we use the notation $\varphi^{\dagger}$ for the conjectures to distinguish them from those defined in Equation 4.1, and we denote by $\Sigma^{\Phi_{\dagger}}$ the induced solution concept.
Fix an economy of length $T$. Recall that 1 implies that no matter what the conjectures are, the set of $\Phi$-solutions for an economy of length $1$ coincide with the set of stable matchings. Therefore, fix a matching through period $T-2$, $m^{T-2}$. For $k\in\pazocal{A}_{T-1}\left(m^{T-2}\right)\cup\pazocal{B}_{T-1}\left(m^{T-2}\right)$, define $k$’s conjectures as follows:
$\varphi^{\dagger}_{T-1}(m^{T-2},k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{T-2},k):\overline{m}_{T}|_{E^{T}_{T}(\overline{m}^{T-1})}\in\mathrm{S}\left(E^{T}_{T}(\overline{m}^{T-1})\right)\right\}.$
Let $\Sigma^{\Phi_{\dagger}}_{T-1}\left(E^{T}_{T-1}\left(m^{T-2}\right)\right)$ denote the set of $\varphi^{\dagger}_{T-1}$-solutions for $E^{T}_{T-1}\left(m^{T-2}\right)$.
Recursively, suppose we have defined $\Sigma^{\Phi_{\dagger}}_{T-t}$ for economies of length $T-t$, and fix a matching through period $t$, $m^{t-1}$. For $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, $k$’s conjectures are given by:
$\varphi^{\dagger}_{t}(m^{t-1},k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1},k):(\overline{m}_{s})_{s=t+1}^{T}|_{E^{T}_{t+1}(\overline{m}^{t})}\in\Sigma^{\Phi_{\dagger}}_{T-t}\left(E^{T}_{t+1}(\overline{m}^{t})\right)\right\},$
and let $\Sigma^{\Phi_{\dagger}}_{T-t+1}\left(E^{T}_{t}(m^{t-1})\right)$ denote the set of solutions given the family $\Phi_{\dagger,\geq t}=\{\varphi^{\dagger}_{s}((m^{t-1},\overline{m}_{t}^{s-1}),k):s\geq t,\overline{m}_{t}^{s-1},k\in\pazocal{A}_{s}(m^{t-1},m_{t}^{s-1})\cup\pazocal{B}_{s}(m^{t-1},m_{t}^{s-1})\}$.
The following statement is an immediate corollary of Theorem 4.1 and the definition of the conjectures $\varphi^{\dagger}$:
@@
For all $T\in\mathbb{N}$, $\Sigma^{\Phi_{\dagger}}_{T}$ is a nonempty-valued correspondence.
A.1 then shows existence issues for solution concepts in dynamic matching markets do not arise solely from the perfection requirement—that is, that the continuation matching also abides by the solution concept—but from restrictions the conjectures place on the period-$t$ matchings.
A.1 extends to dynamic matching markets with irreversible matching the existence result for repeated matching in Kotowski (2019). Indeed, conjectures in Kotowski (2019) consist of all matchings whose continuations are solutions in the continuation economy.
A.1 Rational expectations and omitted details from 4.1
In this section, we formally define the rational expectations solution, discussed at the end of Section 4. We then provide supporting details for the derivation in 4.1.
Rational expectations solution
The rational expectations solution is defined via a double recursion on the length of the economy and the number of agents available to match in the first period. Below, we proceed in three steps. First, we introduce a piece of notation that simplifies the exposition. Second, we define the rational expectations solution for an economy of length $1$, which gives us the initial step in the recursion on the length of the economy. Third, we define the rational expectations solution for a length-$T$ economy.
Notation Fix a period-$t$ matching $m_{t}$ defined over $\overline{A}_{t}\cup\overline{B}_{t}$ and let $k\in\overline{A}_{t}\cup\overline{B}_{t}$ be such that $m_{t}(k)=k$. Define $m_{t}\setminus k$ to be the period-$t$ matching with domain and codomain $\overline{A}_{t}\cup\overline{B}_{t}\setminus\{k\}$ such that for all $k^{\prime}\neq k$, $(m_{t}\setminus k)(k^{\prime})=m_{t}(k^{\prime})$.
Economies of length $T=1$ We first note for economies of length $1$, the rational expectations solution corresponds to the set of stable matchings (4). That is, for all $E^{1}=(A_{1},B_{1})$, $\mathrm{RE}_{1}(E^{1})=\mathrm{S}(E^{1})$.
Economies of length $T\geq 2$ Fix an economy $E^{T}$ and suppose we have defined the rational expectations solution for all economies of length $t\in\{1,\dots,T-1\}$.
Given any set $C\subseteq A_{1}\cup B_{1}$, let $E^{T}_{\setminus C}$ denote the length-$T$ economy in which the agents in $C$ arrive in $t=2$, rather than in $t=1$; that is,
$E^{T}_{\setminus C}=(A_{1}\setminus C,B_{1}\setminus C,A_{2}\cup(A_{1}\cap C),B_{2}\cup(B_{1}\cap C),\dots,A_{T},B_{T}).$
Let $N$ denote the cardinality of $A_{1}\cup B_{1}$. To define the rational expectations solution for $E^{T}$, we need to define the rational expectations solution for $E^{T}_{\setminus C}$ for $C\subseteq A_{1}\cup B_{1}$ for $C$ of cardinality $1,\dots,N$, starting with coalitions of size $N$:
  1. 1.
    When $C$ has cardinality $N$, the rational expectations solution of $E^{T}_{\setminus C}$, $\mathrm{RE}_{T}(E^{T}_{\setminus C})$ corresponds to the rational expectations solution of the induced length $T-1$ economy, $\mathrm{RE}_{T-1}(E^{T}_{\setminus C})$. Formally,
    $\mathrm{RE}_{T}(E^{T}_{\setminus C})=\{\overline{m}\in\pazocal{M}_{T}(E^{T}_{\setminus C}):(\overline{m}_{s})_{s=2}^{T}\in\mathrm{RE}_{T-1}(E^{T}_{\setminus C})\}.$
  2. 2.
    Suppose $C$ has cardinality $m\in\{1,\dots,N-1\}$ and $\mathrm{RE}_{T}$ has been defined for $E^{T}_{\setminus C^{\prime}}$ for all $C^{\prime}:C\subseteq C^{\prime}$. For each $k\in A_{1}\cup B_{1}\setminus C$, define
    $\varphi_{\mathrm{RE},1}(E^{T}_{\setminus C},k)=\{\overline{m}\in\pazocal{M}_{T}(E^{T}_{\setminus C},k):(\overline{m}_{1}\setminus k,(\overline{m}_{s})_{s=2}^{T})\in\mathrm{RE}_{T}(E^{T}_{\setminus(C\cup\{k\})})\}.$
    The rational expectations solution for $E^{T}_{\setminus C}$, $\mathrm{RE}_{T}(E^{T}_{\setminus C})$, is defined as the subset of $\pazocal{M}_{T}(E^{T}_{\setminus C})$ that satisfies 7 when period-1 conjectures are given by $\{\varphi_{\mathrm{RE},1}(E^{T}_{\setminus C},k):k\in A_{1}\cup B_{1}\setminus C\}$ and continuation matchings are elements of $\mathrm{RE}_{T-1}$.
@@
In what follows, we provide supporting details for the derivation in 4.1. First, note should no one match in period $1$, a unique stable matching exists in $t=2$, which leads to the matching:
@@
Second, consider $E_{\setminus\{b_{1}\}}^{2}$. We argue $m_{b_{1}}$ is a rational expectations solution. To illustrate, note $b_{2}$ does not wish to block with $a_{1}$, and $a_{3}$ cannot do better than by matching in $t=1$ with $b_{2}$. The only possibility is that $b_{2}$ waits to be matched. However, $m_{\emptyset}$ is a valid conjecture for $b_{2}$ in $E_{\setminus\{b_{1},b_{2}\}}^{2}$, and $b_{2}$ prefers $a_{3}$ to matching with $a_{1}$ in $t=2$.
Third, consider $E_{\setminus\{a_{3},b_{1}\}}^{2}$. We show $b_{1}$ is always matched to $a_{4}$ in any rational expectations solution, and hence, $b_{1}$ cannot match with $a_{1}$ in a rational expectations solution for $E_{\setminus\{a_{3}\}}^{2}$. In $E_{\setminus\{a_{3},b_{1}\}}^{2}$, three possibilities exist for a matching that is a rational expectations solution: (i) no one matches in $t=1$, which would lead to $m_{\emptyset}$ (though note $m_{\emptyset}$ is blocked by $(a_{1},b_{2})$ in $E_{\setminus\{a_{3},b_{1}\}}^{2}$), (ii) $a_{1}$ matches with $b_{2}$, and (iii) $a_{2}$ matches with $b_{2}$. Under (ii), note the outcome of side-$A$ proposing deferred acceptance in $t=2$ among $a_{2},a_{3},a_{4},b_{1},b_{3},b_{4}$ is
@@
so any rational expectations solution for $E_{\setminus\{a_{3},b_{1}\}}^{2}$ that matches $a_{1},b_{2}$ in $t=1$ (if any) gives $b_{1}$ at least the payoff of matching with $a_{4}$. Similarly, under (iii), note the outcome of side-$A$ proposing deferred acceptance in $t=2$ among $a_{1},a_{3},a_{4},b_{1},b_{3},b_{4}$ is
@@
so that any rational expectations solution for $E_{\setminus\{a_{3},b_{1}\}}^{2}$ that matches $a_{2},b_{2}$ in $t=1$ (if any) gives $b_{1}$ at least the payoff of matching with $a_{4}$. It follows that in $E_{\setminus\{a_{3}\}}^{2}$, $b_{1}$ has to match with $a_{4}$.
Fourth, consider $E_{\setminus\{a_{2}\}}^{2}$. We argue $m_{a_{2}}$ is a rational expectations solution for that economy. To do so, we need to argue $b_{1}$ cannot block. Consider then the economy $E_{\setminus\{a_{2},b_{1}\}}^{2}$ and note $m_{b_{1}}$ is a rational expectations solution for the same reason it is a rational expectations solution in $E_{\setminus\{b_{1}\}}^{2}$. Thus, by blocking $m_{a_{2}}$, $b_{1}$ cannot expect to get more than the payoff of matching with $a_{1}$ in $t=2$. It follows that $m_{a_{2}}$ is a rational expectations solution for $E_{\setminus\{a_{2}\}}^{2}$.
Finally, we argue $m^{\star}$ in the example is a candidate matching as defined in Section 4. Note that because of $m_{a_{2}}$, $a_{2}$ is willing to match with $b_{2}$ in the one-period economy induced by the conjectures, and similarly, $b_{1}$ is willing to match with $a_{1}$. Because $a_{2}$ and $b_{2}$ top-rank each other in the one-period economy induced by the conjectures, they match together in any stable matching, which then justifies $a_{1}$ and $b_{1}$ matching, and $a_{3}$ remaining unmatched. When $\{a_{3},a_{4}\}$ are the $t=2$ agents on side $A$, a unique stable matching exists, which is specified by $m^{\star}_{2}$.